OpenStudy (anonymous):

what value is solution of the equation 3-2x=-5 A.2 B.4 C.-2 D.1/5

OpenStudy (anonymous):

the answer is b

OpenStudy (anonymous):

3+5=2x 8=2x x=8/2 x=4 So, its B :)